# InstructBioMol: Advancing Biomolecule Understanding and Design Following Human Instructions

### v1

@@ Methods
InstructBioMol: Advancing Biomolecule Understanding and Design Following Human Instructions
@@ 1.2 Training Details
During continual pretraining, we set batch size to 32 and fix sequence length to 512 tokens. To ensure balanced training on molecule, protein, and natural language data, the sampling ratio for these three types of data is fixed at 1:1:1. In instruction-tuning, we use a batch size of 24, with the maximum sequence length set to 512. By setting hyperparameters, we control the sampling ratio of different types of instruction data. The sampling ratios for the datasets used in stage-1 and stage-2 of instruction-tuning are detailed in Table 1.
@@ 1.3 Inference Settings
For inference on downstream tasks, we load the model using the bfloat16 data format. The specific inference hyperparameters for each task are detailed in Table 2.
@@ 2 Baselines
To validate the effectiveness of InstructBioMol, we conduct experiments comparing it with various baseline models. In molecule captioning and description-based molecule generation tasks, the selected baseline models include MolT5 [14], BioT5 [16], BioT5+ [18], MolReGPT [31], InstructMol [33] and ChemDFM [32]. For comparisons on other tasks, baselines are categorized into two groups. The first group comprises pre-trained models, including Mol-Instructions [17], InstructProtein [15], ProtT3 [34], BioMedGPT [19], ProteinDT [35], DrugGPT [42] and RFdiffusionAA [5]. These models are all downloaded from their official repositories and evaluated on the test set. The second group consists of baseline models constructed by us using the general-purpose Large Language Model GPT-3.5. It includes three variants: zero-shot, 5-shot random, and 5-shot similarity. In the zero-shot setting, we directly pose task-specific questions to the GPT-3.5 model. In the 5-shot random setting, five examples from the training set are randomly selected as in-context demonstrations for each test entry. In the 5-shot similarity setting, the in-context learning paradigm is also adopted, but the demonstrations are required to be the five most similar examples from the training set relative to the query. The method for computing similarity depends on the data type of the input query: when query is in natural language, TF-IDF with cosine similarity is used as the text similarity measure; for protein sequence queries, MMseq2 [76] is employed to calculate protein similarity; and for molecule queries, molecular fingerprint similarity [29, 72] is used. The specific input format for the employed GPT baseline is detailed in Table 3.
@@ 3.1 Examples of Datasets
We provide the example entries of continual pretraining dataset in Table 4, and the example entries of instruction-tuning dataset in Table 5, Table 6 and Table 7.
@@ 3.2 Split of Datasets
To evaluate the model’s performance, we divide the dataset used for the stage-2 instruction tuning into training and test sets. For the dataset aligning molecules with natural language, we adopt the data split from ref. 14. For the dataset aligning proteins with natural language, we use the training data defined in ref. 19 and randomly select 3000 samples from the test set as our evaluation data. Statistics of the above two datasets are in Table 8. For the dataset aligning molecules with proteins, we account for the specific nature of the tasks. In the task of generating drug-like molecules for proteins, a single protein typically corresponds to multiple molecules. Conversely, in the task of generating enzymes for substrate molecules, a single substrate often corresponds to multiple proteins. Accordingly, we split the dataset as follows: for the former task, we select 100 target proteins and their corresponding molecules as the test set. For the latter task, we select 100 target substrates and their corresponding proteins as the test set. The specific sizes of each dataset split are detailed in Table 9.
4 Additional Results
@@ 4.1 Case Analysis for Experimental Results of Understanding and Designing Molecules
For the task of description-based molecule generation, several case examples of the ground truth and generation are presented in Figure 1. Overall, InstructBioMol demonstrates a relatively accurate analysis of molecular structure, function, origin, etc. In the task of description-based molecule generation, InstructBioMol is capable of generating molecules that are completely consistent with the ground truth in certain cases, such as molecules PubChem-CID-5281294 and PubChem-CID-31284 shown in Figure 2. This demonstrates the strong molecular design capability of InstructBioMol. However, in some other cases, the molecules generated by InstructBioMol show some discrepancies with the ground truth. Our analysis suggests that this may be due to inadequate handling of certain functional groups. For example, for PubChem-CID-118429016 and PubChem-CID-123953, the model omits certain functional groups (a hydroxyl group and a phosphate group, respectively). For PubChem-CID-179394, the model generates a chemically atypical P(O)(O)(O)O group. Overall, InstructBioMol exhibits a high level of accuracy in molecule generation tasks and shows potential for applications in fields such as drug discovery, and further optimization may be required in practical applications.
@@ 4.2 Case Analysis for Experimental Results of Understanding and Designing Proteins
In the protein function answering task, InstructBioMol generates results that closely resemble the ground truth for certain cases, such as Q9NRY2 and P73070 in Table 10. Furthermore, we observe that in some cases, the generated descriptions tend to be more detailed. For example, in the case of Q9Y2G3, the generated description includes detailed information on vesicle formation, lipid signal molecule uptake, and the establishment of the thrombopoietin gradient in platelets. Similarly, for Q9FY89, the generated description provides a more comprehensive explanation of the formation of multivesicular bodies (MVBs), specifically describing the formation mechanism of intraluminal vesicles (ILVs) within MVBs, including the invagination and scission of the endosomal membrane. It further elaborates on the function of MVBs, such as transporting their contents to lysosomes for the degradation of membrane proteins, receptors, lysosomal enzymes, and lipids. For P0CP67, the generated description offers more detailed and in-depth functional information, identifying the specific targets of the protein’s action (e.g., components of AP-1, c-Jun, and ATF2) and potential biological processes involved (e.g., regulation of circadian clock). Although InstructBioMol may provide researchers with deeper insights, further experimental validation is necessary to confirm these findings.
For the task of text-based protein generation, we present two examples in Figure 3. For the ground truth proteins, we utilize the protein structures predicted by AlphaFold Protein Structure Database [69]. For the generated protein sequences, we predict their structures using ColabFold [77]. Besides sequence similarity metrics Identity, Alignment, and BLOSUM Substitution, we also compare structural similarity metrics: TM-Score [78] and LDDT [79]. The results demonstrate that InstructBioMol is capable of de-novo design of proteins, with the designed proteins exhibiting a high degree of structural similarity to ground truth. This suggests that InstructBioMol holds significant potential in designing proteins tailored to specific functional descriptions, acting as an effective copilot to assist researchers in protein design.
Table 1: Sampling ratios of different types of instruction data during instruction-tuning. Note that in practice, the sampling ratios are scaled proportionally to ensure that the sum of the ratios for all data equals 1.
@@
Table 2: Inference hyperparameters for downstream tasks.
@@
Table 3: In-context learning examples of the GPT baseline across different tasks, Where XX represents data-specific natural language descriptions, protein sequences, or molecular sequences. In the few-shot setting, the input to the GPT model consists of a template, in-context demonstrations, and a question; in the zero-shot setting, the input consists of a template and a question.
@@
Table 4: Examples of the continual pertaining data.
@@
Table 5: Examples of the instruction-tuning data on molecule-natural language alignment.
@@
Table 6: Examples of the instruction-tuning data on protein-natural language alignment.
@@
Table 7: Examples of the instruction-tuning data on molecule-protein alignment.
@@
Table 8: Statistics of datasets for molecule-natural language alignment and protein-natural language alignment.
@@
Table 9: Statistics of datasets for molecule-protein alignment. # of Entries and # of Targets denote the number of data entries and the number of targets, respectively.

### v2

@@ Acknowledgements
Advancing Biomolecular Understanding and Design Following Human Instructions
@@ 1 Implementation of InstructBioMol
1.1 Details of Multimodal Encoders
The molecule 2D-graph encoder, $f_{m}^{2D}$, is a pre-trained 5-layer GIN [42, 43] with an output dimension of $d_{m}^{2D}=300$. The aggregation function of GIN can be formulated as $\mathbf{h}_{v}^{k}=\mathrm{MLP}^{k}((1+\epsilon^{k})\mathbf{h}_{v}^{k-1}+\sum_{u\in\mathcal{N}_{(v)}}\mathbf{h}_{u}^{k-1})$, where $\mathbf{h}_{v}^{k}$ is the embedding of node $v$ at level $k$, and $\mathcal{N}_{(v)}$ represents the neighbors of node $v$. The molecule 3D-structure encoder, $f_{m}^{3D}$, is Geoformer [44]. It is a novel geometric Transformer that integrates Interatomic Positional Encoding (IPE) to effectively model molecular structures, with an output dimension of $d_{m}^{3D}=512$. For protein encoding, we use pre-trained ESM2-35M [41] and SaProt-35M [45] as the protein 1D-sequence encoder $f_{p}^{1D}$ and 3D-structure encoder $f_{p}^{3D}$, respectively, with an output dimension of 480 for both.
@@ 1.3 Training Details
During continual pretraining, we set batch size to 32 and fix sequence length to 512 tokens. To ensure balanced training on molecule, protein, and natural language data, the sampling ratio for these three types of data is fixed at 1:1:1. In instruction-tuning, we use a batch size of 24, with the maximum sequence length set to 512. By setting hyperparameters, we control the sampling ratio of different types of instruction data. The sampling ratios for the datasets used in stage-1 and stage-2 of instruction-tuning are detailed in Supplementary Table 1.
The downstream task “designing protein and molecule simultaneously from descriptions” is structured as follows: $(x_{instruction},x_{text})\to(y_{prot},y_{mol})$, where $x_{instruction}$ is "Please generate the corresponding protein sequence based on the following description of the protein and generate a molecule that can interact with this protein". $x_{text}$ is the natural language description of the protein, and $y_{prot}$, $y_{mol}$ are generated protein sequence and molecule, respectively. During training, we perform LoRA [38] fine-tuning (with hyperparameters set to $r_{lora}=32,\alpha_{lora}=32$) on InstructBioMol. LoRA fine-tuning is highly efficient, updating only about 0.5% of the model parameters.
@@ 1.4 Inference Settings
For inference on downstream tasks, we load the model using the bfloat16 data format. The specific inference hyperparameters for each task are detailed in Supplementary Table 2. In the experiments of protein understanding and design following human instructions, the experiments are repeated with three different random seeds.
@@ 2 Baselines
In molecule captioning and description-based molecule generation tasks, the selected baselines include MolT5 [15], BioT5 [17], BioT5+ [19], MolReGPT [27], InstructMol [29] and ChemDFM [28]. InstructMol is incapable of molecule generation.
For protein understanding and design, target-protein-based drug design, and substrate-based enzyme design tasks, baselines are categorized into two groups. The first group comprises pre-trained models, including Mol-Instructions [18], InstructProtein [16], ProtT3 [30], BioMedGPT [20], ProteinDT [31], Pocket2Mol [33], AR-SBDD [34], TargetDiff [35], DrugGPT [36] and RFdiffusionAA [5]. These models are all downloaded from their official repositories and evaluated on the test set. In protein understanding and design task, BioMedGPT and ProtT3 cannot generate protein sequences, while ProteinDT cannot answer protein property-related questions using natural language. We also train a variant of ProteinDT that incorporates motif information (ProteinDT+motif) on description-based protein design task. In this model, we add motif embeddings to the protein encoder during the first stage of training, which aligns text and protein pairs. Specifically, the protein representation is generated as: $p_{p}\circ[f_{p}(x_{p})\oplus\mathbf{x}_{motif}]$, where $p_{p}$ is a projector, $f_{p}$ is the protein encoder, $x_{p}$ is the protein sequence, and $\mathbf{x}_{motif}$ is the motif embedding. Subsequently, we continue training the protein facilitator and protein decoder following the original procedure. The second group consists of baseline models constructed by us using the general-purpose Large Language Model GPT-3.5. It includes three variants: zero-shot, 5-shot random, and 5-shot similarity. In the zero-shot setting, we directly pose task-specific questions to the GPT-3.5 model. In the 5-shot random setting, five examples from the training set are randomly selected as in-context demonstrations for each test entry. In the 5-shot similarity setting, the in-context learning paradigm is also adopted, but the demonstrations are required to be the five most similar examples from the training set relative to the query. The method for computing similarity depends on the data type of the input query: when query is in natural language, TF-IDF with cosine similarity is used as the text similarity measure; for protein sequence queries, MMseq2 [85] is employed to calculate protein similarity; and for molecule queries, molecular fingerprint similarity [49, 74] is used. The specific input format for the employed GPT baseline is detailed in Supplementary Table 3.
For designing protein and molecule simultaneously from descriptions, baselines consist of three groups. (1) The first group is GPT-3.5, which includes two variants: GPT-3.5 (5-shot, random) and GPT-3.5 (5-shot, similarity). These models utilize an in-context learning paradigm. The former randomly selects 5 examples from the training set as prompts, while the latter selects the 5 most similar examples to the given query. (2) The second group consists of fine-tuned Large Language Models. This includes Llama [7] , Mol-Instructions-molecule, Mol-Instructions-protein, BioT5, and BioT5+. Mol-Instructions-molecule and Mol-Instructions-protein are Llama-based models trained with molecule-text and protein-text instructions from the Mol-Instructions dataset [18], respectively. All models in this group are fine-tuned on the training set. For Llama, Mol-Instructions-molecule, and Mol-Instructions-protein, we use LoRA fine-tuning, which aligns with the approach used in InstructBioMol. For BioT5 and BioT5+, we fine-tune all parameters. The number of fine-tuning steps is kept consistent across all models. (3) The third group is a sequential ensemble approach: GPT-3.5 (5-shot, similarity) + DrugGPT. This method combines the most competitive techniques for generating proteins from natural language (GPT-3.5 (5-shot, similarity)) and generating molecules from proteins (DrugGPT). First, GPT-3.5 (5-shot, similarity) is used to generate proteins from text, followed by the use of DrugGPT to generate molecules based on the proteins.
@@ 3.1 Examples of Datasets
We provide the example entries of continual pretraining dataset in Supplementary Table 4, and the example entries of instruction-tuning dataset in Supplementary Table 5, Supplementary Table 6 and Supplementary Table 7.
@@ 3.2 Split of Datasets
To evaluate the model’s performance, we divide the dataset used for the stage-2 instruction tuning into training and test sets. For the dataset aligning molecules with natural language, we adopt the data split from ref. 15. For the dataset aligning proteins with natural language, we use the training data defined in ref. 20 and randomly select 3000 samples from the test set as our evaluation data. Statistics of the above two datasets are in Supplementary Table 8. For the dataset aligning molecules with proteins, we account for the specific nature of the tasks. In the task of generating drug-like molecules for proteins, a single protein typically corresponds to multiple molecules. Conversely, in the task of generating enzymes for substrate molecules, a single substrate often corresponds to multiple proteins. Accordingly, we split the dataset as follows: for the former task, we select 100 target proteins and their corresponding molecules as the test set. For the latter task, we select 100 target substrates and their corresponding proteins as the test set. The specific sizes of each dataset split are detailed in Supplementary Table 9. For designing protein and molecule simultaneously from descriptions task, we curate triplets from the BindingDB and SwissProt_Description datasets, each consisting of a protein description, its corresponding protein amino acid sequence, and a binding molecule, aligned by the UniprotID. To ensure robust evaluation, we implement data filtering during the training-test split to avoid data leakage between the training and test sets. Our final training set contains 334,216 samples, with 100 protein descriptions in the test set.
4 Additional Experiments
4.1 Molecule Editing
we performed LoRA fine-tuning (with hyperparameters $r_{lora}=32,\alpha_{lora}=32$) on the pretrained InstructBioMol model after the instruction-tuning stage. Specifically, we followed the experimental setup of ChatDrug [21] and conducted experiments on both single-target molecule editing tasks (Task IDs 101-108) and multi-target molecule editing tasks (Task IDs 201-206). To construct the training dataset for these tasks, we randomly selected 1,000 molecules from the retrieval database provided by ChatDrug as the original molecules. ChatDrug was then used to generate pairs of original and successfully edited molecules for each task. We applied LoRA to perform multi-task fine-tuning on InstructBioMol, covering 8 single-target editing tasks and 6 multi-target editing tasks. The task format is $(x_{instruction},x_{multimodal\_mol})\to y_{mol}$, with the instructions for different tasks provided in Supplementary Table 10.
It is worth noting that LoRA fine-tuning is a highly efficient method, as the model only needs to optimize approximately 0.5% of the total parameters. This allows the training to be completed in just six hours on NVIDIA 3090 GPU (24GB memory). After fine-tuning, we obtained a model capable of handling various molecular editing tasks. The model was then evaluated on the test set provided by ChatDrug, using five random seeds for repeated experiments. The hit ratio (the proportion of successful edits) was used as the evaluation metric.
In Supplementary Tables 11 and 12, we report the experimental results for single-target and multi-target editing tasks, respectively. Each editing task includes two different threshold settings: a loose threshold and a strict threshold, resulting in a total of 28 tasks. The average performance of the model across all tasks is also provided. It is evident that InstructBioMol demonstrates the best overall performance. In both single-target and multi-target editing tasks, InstructBioMol achieves the highest average performance and outperforms the other models in more than half (15/28) of the tasks. More importantly, we also observe that InstructBioMol shows particularly significant performance improvement in the more challenging multi-target molecule editing tasks, with a 17.8% increase. This result underscores InstructBioMol’s ability not only to improve editing accuracy but also to enhance efficiency when handling complex molecule editing tasks, further showcasing its strengths in the molecular editing domain. We have also updated the code repository with the training and inference code for LoRA fine-tuning on molecular editing tasks, along with the associated data and model weights.
Through this molecule editing experiment, we demonstrate that InstructBioMol can efficiently adapt to various downstream tasks via LoRA fine-tuning. This flexibility arises from InstructBioMol being a unified natural language-molecule-protein model, trained on large-scale and diverse datasets, enabling excellent any-to-any alignment. As a result, InstructBioMol can tackle a wide range of complex tasks in the biomolecular domain, providing a strong foundation for further research and applications. This highlights InstructBioMol’s tremendous potential as a foundational model for biomolecular tasks.
4.2 Designing Molecules based on Pairs of Target Proteins
Generalizing to multiple molecule-protein alignments is indeed an exciting area of exploration. As mentioned in the Discussion Section, due to resource constraints, we did not include all potential biomolecular tasks in the current study. However, the model architecture and training framework we propose are designed to be highly generalizable and scalable, allowing for easy adaptation to a variety of tasks.
To explore how the model could generalize to multiple molecule-protein alignment tasks—such as generating molecules that bind or do not bind to several proteins—we found that LoRA fine-tuning offers a simple yet highly effective solution. LoRA only requires optimizing 0.5% of the model parameters, and it can be completed with high computational efficiency on 24GB NVIDIA 3090 GPU. Specifically, we applied LoRA fine-tuning to conduct experiments on generating molecules that bind to a pair of proteins. First, we collected data on protein pairs and molecules that bind to both of them simultaneously, and divided the data into training and test sets. The training set contains 102,075 protein pairs-molecule binding triplets, and the test set includes 30 pairs of protein targets. We adopted the same training paradigm as in the instruction-tuning phase: $(x_{instruction},x_{multimodal\_prot1},x_{multimodal\_prot2})\to y_{mol}$, where $x_{instruction}$ is “Generate a drug molecule that binds to these target proteins", $x_{multimodal\_prot1}$ and $x_{multimodal\_prot2}$ are the two target proteins, and $y_{mol}$ is the generated molecule. The hyperparameters for LoRA were set to $r=32$ and $\alpha=32$, and the fine-tuning process required only about half an hour.
In the experiment, we used Vina Score, High Affinity, and Success Rate as evaluation metrics. For calculating High Affinity and Success Rate, we required that the generated molecules satisfy the conditions on both target proteins. Since this is a very novel task and no specialized models have been designed for it, we compared InstructBioMol with the general model GPT-3.5 and its in-context learning variant. The experimental results in Supplementary Figure 2 show that InstructBioMol significantly outperforms the baseline models. This highlights the model’s ability to quickly adapt to new tasks through LoRA-based instruction-tuning, showcasing its excellent scalability and generalization. Our findings suggest that InstructBioMol’s any-to-any alignment allows it to handle complex tasks, further underscoring its potential for broad applications in biomolecular research.
4.3 Ablation Analysis
Analysis of Training Strategy.
We analyze the impact of different training strategies. Supplementary Figure 3a compares InstructBioMol with two variants: one that retains continual pretraining and stage-2 instruction-tuning but removes stage-1 instruction-tuning ("w/ continual-pretraining, w/o stage-1 instruction-tuning"), and another that only retains stage-2 instruction-tuning, removing both continual pretraining and stage-1 instruction-tuning ("w/o continual-pretraining, w/o stage-1 instruction-tuning"). Our findings are as follows:
For tasks related to molecule understanding and design (e.g., molecule captioning and description-based generation), continual pretraining has a minor impact, while stage-1 instruction-tuning is more crucial. In molecule captioning, removing stage-1 instruction-tuning results in a 7.3% performance drop, with an additional 1.2% decrease after removing continual pretraining. In molecule generation, removing stage-1 instruction-tuning leads to a 12.4% drop in accuracy, with only a 1.1% further decrease after removing continual pretraining. We attribute this to the role of molecular IUPAC name data in stage-1 instruction-tuning, which aligns molecular structure with natural language. For description-based protein generation, continual pretraining plays a more significant role, while stage-1 instruction-tuning has a smaller impact. Removing stage-1 instruction-tuning decreases the Identity by 3.6%, while further removal of continual pretraining results in an 8.9% drop. This suggests that the evolutionary information captured during continual pretraining, particularly from broad protein sequence datasets, enhances the quality of generated proteins.
Analysis of Multimodal Data Integration.
Next, we explore the impact of incorporating multimodal data on model performance. As shown in Supplementary Figure 3b, we compare InstructBioMol and its variants on molecule captioning and protein function answering tasks. Variants include the removal of the 2D encoder (w/o 2D encoder) and 3D encoder (w/o 3D encoder) for molecular data, as well as the removal of the 1D encoder (w/o 1D encoder) and 3D encoder (w/o 3D encoder) for protein data. Additionally, we consider the removal of motif prompts (w/o motif) and the entire Motif-Guided Multimodal Feature Extraction Module (w/o multimodal). InstructBioMol and its variants here exclude both continual pretraining and stage-1 instruction-tuning.
The results reveal that removing the Motif-Guided Multimodal Feature Extraction Module significantly decreases performance, with average drops of 4.7% and 7.1% on molecular and protein tasks, respectively. This highlights the importance of multimodal feature extraction, which integrates information from different modalities to provide more comprehensive and accurate representations. Further analysis shows that 3D data has minimal impact on molecular tasks, but is more crucial for protein tasks, where the removal of the 3D encoder leads to a 4.2% decline. This suggests that 3D structures are essential for identifying protein functions. Overall, multimodal data enhances performance in biomolecular tasks, emphasizing its importance in these domains.
5 Additional Analysis
5.1 Analysis of Motifs in Biomolecules
In Supplementary Figures 4 and 5, we provide examples of shared motifs in both molecules and proteins, along with their corresponding textual descriptions. By examining these examples, we observe a clear pattern: when molecules or proteins share common motifs, their associated natural language descriptions also show notable similarities. This observation suggests a strong correlation between motif information and natural language descriptions, highlighting the critical role motifs play in capturing the functional properties of biomolecules. As such, the incorporation of motif-based features into the model improves its ability to capture these similarities and better understand molecular functions, ultimately enhancing model performance.
5.2 Novelty of the Generated Biomolecules
We compared two aspects of novelty [86]: (1) the proportion of generated drugs or enzymes that do not appear in the reference set, and (2) the proportion that does not appear in the training set. The results are shown in Supplementary Table 13 and Supplementary Table 14. In the drug generation task, our model achieved 100% novelty with respect to the reference set, meaning that all the generated molecules were novel and did not appear in the reference data. This highlights the ability of our approach to expand the known drug space for the given targets. Additionally, when compared to the training set, our method reached a 92.9% novelty rate, representing a 1.5% improvement over the baseline method, DrugGPT. In the enzyme generation task, our model achieved novelty scores of 99.8% compared to the reference set, and 98.0% compared to the training set. While these novelty scores are slightly lower than some baseline methods, they still surpass GPT-3.5 (5-shot, similarity). More importantly, the enzymes generated by our model show substantial improvements in substrate-enzyme interaction, as illustrated in Figure 3d of the Manuscript. These results underscore the potential of our approach to discover novel drugs and enzymes that have not been previously reported or recorded, emphasizing its capability to contribute to the exploration of new biomolecular spaces.
@@ 5.3 Case Analysis for Experimental Results of Understanding and Designing Molecules
For the task of description-based molecule generation, several case examples of the ground truth and generation are presented in Supplementary Figure 6. Overall, InstructBioMol demonstrates a relatively accurate analysis of molecular structure, function, origin, etc. In the task of description-based molecule generation, InstructBioMol is capable of generating molecules that are completely consistent with the ground truth in certain cases, such as molecules PubChem-CID-5281294 and PubChem-CID-31284 shown in Supplementary Figure 7. This demonstrates the strong molecular design capability of InstructBioMol. However, in some other cases, the molecules generated by InstructBioMol show some discrepancies with the ground truth. Our analysis suggests that this may be due to inadequate handling of certain functional groups. For example, for PubChem-CID-118429016 and PubChem-CID-123953, the model omits certain functional groups (a hydroxyl group and a phosphate group, respectively). For PubChem-CID-179394, the model generates a chemically atypical P(O)(O)(O)O group. Overall, InstructBioMol exhibits a high level of accuracy in molecule generation tasks and shows potential for applications in fields such as drug discovery, and further optimization may be required in practical applications.
@@ 5.4 Case Analysis for Experimental Results of Understanding and Designing Proteins
In the protein function answering task, InstructBioMol generates results that closely resemble the ground truth for certain cases, such as Q9NRY2 and P73070 in Supplementary Table 15. Furthermore, we observe that in some cases, the generated descriptions tend to be more detailed. For example, in the case of Q9Y2G3, the generated description includes detailed information on vesicle formation, lipid signal molecule uptake, and the establishment of the thrombopoietin gradient in platelets. Similarly, for Q9FY89, the generated description provides a more comprehensive explanation of the formation of multivesicular bodies (MVBs), specifically describing the formation mechanism of intraluminal vesicles (ILVs) within MVBs, including the invagination and scission of the endosomal membrane. It further elaborates on the function of MVBs, such as transporting their contents to lysosomes for the degradation of membrane proteins, receptors, lysosomal enzymes, and lipids. For P0CP67, the generated description offers more detailed and in-depth functional information, identifying the specific targets of the protein’s action (e.g., components of AP-1, c-Jun, and ATF2) and potential biological processes involved (e.g., regulation of circadian clock). Although InstructBioMol may provide researchers with deeper insights, further experimental validation is necessary to confirm these findings.
For the task of text-based protein generation, we present two examples in Supplementary Figure 8. For the ground truth proteins, we utilize the protein structures predicted by AlphaFold Protein Structure Database [65]. For the generated protein sequences, we predict their structures using ColabFold [87]. Besides sequence similarity metrics Identity, Alignment, and BLOSUM Substitution, we also compare structural similarity metrics: TM-Score [88] and LDDT [89]. The results demonstrate that InstructBioMol is capable of de-novo design of proteins, with the designed proteins exhibiting a high degree of structural similarity to ground truth. This suggests that InstructBioMol holds significant potential in designing proteins tailored to specific functional descriptions, acting as an effective copilot to assist researchers in protein design.
Supplementary Table 1: Sampling ratios of different types of instruction data during instruction-tuning. Note that in practice, the sampling ratios are scaled proportionally to ensure that the sum of the ratios for all data equals 1.
@@
Supplementary Table 2: Inference hyperparameters for downstream tasks.
@@
Designing protein and molecule from descriptions $top\_p=0.7,temperature=1$
Supplementary Table 3: In-context learning examples of the GPT baseline across different tasks, Where XX represents data-specific natural language descriptions, protein sequences, or molecular sequences. In the few-shot setting, the input to the GPT model consists of a template, in-context demonstrations, and a question; in the zero-shot setting, the input consists of a template and a question.
@@
Designing protein and molecule simultaneously from descriptions
You are a biologist. Given the protein description, your task is to design a new protein and a molecule that can bind to this protein using your experienced knowledge. Please respond only with the amino acid sequence using a sequence of the capitalized initial letters of 20 amino acids and SMILES representation of molecule, separated by ‘,’, and DO NOT reply with others. Please strictly follow the format, no other information can be provided. Protein Description: XX; Protein and Molecule: XX, XX . … Protein Description: XX; Protein and Molecule: XX, XX . Protein Description: XX; Protein and Molecule:
Supplementary Table 4: Examples of the continual pertaining data.
@@
Supplementary Table 5: Examples of the instruction-tuning data on molecule-natural language alignment.
@@
Supplementary Table 6: Examples of the instruction-tuning data on protein-natural language alignment.
@@
Supplementary Table 7: Examples of the instruction-tuning data on molecule-protein alignment.
@@
Supplementary Table 8: Statistics of datasets for molecule-natural language alignment and protein-natural language alignment.
@@
Supplementary Table 9: Statistics of datasets for molecule-protein alignment. # of Entries and # of Targets denote the number of data entries and the number of targets, respectively.